Goodee Bakery is considering a change in its inventory valuation method. Goodee Bakery currently uses the FIFO method and is considering a change to the LIFO method. Goodee Bakery started the year on January 1 with inventory at a FIFO cost of $23,500 and a LIFO cost of $21,000. The ending inventory on December 31 is $25,600 at FIFO cost and $21,300 at LIFO cost. The LIFO effect is OA, OB. OC. O D. $2,500 $4,300 $1,800 $6,800
Answer #1

Lifo reserve on December 31 = 25,600 - 21,300

= 4,300

Lifo reserve on January 1 = 23,500 - 21,000

= 2,500

Lifo effect = 4,300 - 2,500

= 1,800

Option C is the answer